Find the value of


C0 + 1*C1 + 2*C2 +.......+ n*Cn.

Grade:12th Pass

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Speed,

let f(x)=y=(1+x)n

y'=C0+1*x*C1+2x2C2+...nxnCn

let x=1,then y=2n,y'=n(1+x)n-1=n(1+1)n-1=n*2n-1

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Speed !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

let f(x)=y=(1+x)n

y'=

Ashwin Muralidharan IIT Madras
12 years ago

Hi Speed Racer,

 

Kindly note, when differentiating, you would get n*(1+x)^n-1.

 

And so the answer would be n2n-1+1.

Wrongly mentioned as n2n+1.

 

Best Regards,

Ashwin (IIT Madras).